### v2

# Ditto: Motion-Space Diffusion for Controllable Realtime Talking Head Synthesis

Abstract
Recent advances in diffusion models have revolutionized audio-driven talking head synthesis. Beyond precise lip synchronization, diffusion-based methods excel in generating subtle expressions and natural head movements that are well-aligned with the audio signal. However, these methods are confronted by slow inference speed, insufficient fine-grained control over facial motions, and occasional visual artifacts largely due to an implicit latent space derived from Variational Auto-Encoders (VAE), which prevent their adoption in realtime interaction applications. To address these issues, we introduce Ditto, a diffusion-based framework that enables controllable realtime talking head synthesis. Our key innovation lies in bridging motion generation and photorealistic neural rendering through an explicit identity-agnostic motion space, replacing conventional VAE representations. This design substantially reduces the complexity of diffusion learning while enabling precise control over the synthesized talking heads. We further propose an inference strategy that jointly optimizes three key components: audio feature extraction, motion generation, and video synthesis. This optimization enables streaming processing, realtime inference, and low first-frame delay, which are the functionalities crucial for interactive applications such as AI assistants. Extensive experimental results demonstrate that Ditto generates compelling talking head videos and substantially outperforms existing methods in both motion control and realtime performance.
footnotetext: † Corresponding Author
Figure 1: The summary of the proposed Ditto, where $\mathcal{T}$ is the DiT for motion generation, $\mathcal{H}$ is the HuBERT for audio feature extraction, $\mathcal{M}$ is the Motion Extractor, $\mathcal{F}$ is the Appearance Feature Extractor and $\mathcal{G}$ is the Face Renderer.
1 Introduction
Audio-driven talking head synthesis [1, 2, 3, 4, 5, 6, 7] has emerged as a prominent research topic in recent years. In particular, one-shot methods [8, 9, 10, 11], which enable video generation from a single photograph, substantially reduce the data prerequisite for target identities, allowing for the creation of talking-head videos without relying on templates. In early research, approaches based on Generative Adversarial Networks (GANs) [12] are the predominant solution for this task. These studies demonstrate relatively accurate lip synchronization capabilities. However, they still lack diversity and realism in terms of speaking styles, emotions, and head movements. Recently, diffusion-based techniques [13, 14, 15], with EMO [3] as a notable example, have made substantial progress. These methods maintain excellent lip synchronization while further enhancing the liveliness and realism of generated videos. They capture subtle facial expressions, synchronize head movements with speech rhythm, and convey lifelike gaze interactions, which are crucial traits to achieve truly convincing synthesized results.
However, two critical issues hinder the broad application of these methods. The first issue is the lack of fine-grained control over facial motion. Existing approaches predominantly rely on audio and other weak conditions to influence the generated output, making it hard to enforce precise and delicate control over subtle aspects of a talking head, such as the driving regions, emotions and head movements. This limitation complicates the timely modification of small flaws in the generated results and the integration with undriven body regions, restricting applications in diverse scenarios. Moreover, small artifacts in teeth and hair are frequently observed in general text-to-image diffusions for portrait generation, further constraining the performance of these talking head approaches. The second issue is the relatively slow inference process. Currently, most diffusion-based approaches can hardly finish inference in realtime on a single GPU, which is a critical requirement for interactive scenarios. Therefore, although these methods offer enhanced vividness and expressiveness, the inference speed constraint largely prevents their deployment in realtime interactive applications such as AI assistants.
Looking into the aforementioned issues, we argue that their root cause is the use of a redundant, implicit, and inexplicable latent space for training the diffusion model. These approaches apply general text-to-image frameworks to the talking-head domain without decoupling the motion generation and photorealistic neural rendering stages [16]. Specifically, most current approaches adopt the Dual U-Net architecture introduced in EMO, training diffusion models within the latent space of a fixed pre-trained VAE. However, this generic latent space entangles motion and appearance representations, complicating both the diffusion process and subsequent decoding. While VASA-1 [17] achieved real-time inference by decoupling motion generation and facial rendering, it relies on an implicit facial representation, precluding explicit control over the generated motion. Nevertheless, this approach demonstrates the feasibility of integrating diffusion models with the representation space of facial reenactment.
We propose Ditto, a diffusion-based framework specifically designed for one-shot talking head synthesis. Ditto can produce lifelike videos from a single portrait image, enabling fine-grained facial motion control while achieving realtime inference. Our core idea is to leverage diffusion solely for identity-agnostic motion generation and employ a specialized one-shot face renderer for rapid, high-quality video synthesis. This design reduces the training complexity of the diffusion module and allows separate optimization of inference processes between modules.
Specifically, we define an identity-agnostic motion space that exclusively encompasses facial and head motions pertinent to talking-head animations. This space has a dimensionality ten times lower than conventional VAE spaces and establishes semantic correspondences with facial motions, thereby enabling fine-grained control over the generated results. We train a Diffusion Transformer (DiT) [18] in this space using audio, emotion labels, and eye states as conditions. Subsequently, we employ a dedicated one-shot decoder to render the generated motion features, achieving rapid and high-quality video synthesis.
Besides, we implement inference optimization suitable for streaming applications across three primary modules: audio feature extraction, DiT inference, and video synthesis. To our knowledge, our approach is the first to achieve low first-frame delay and realtime streaming inference for diffusion-based talking head generation. Our approach combines lifelike synthesis with rapid generation, providing foundational capabilities for innovative realtime dialogue applications such as AI assistants.
2 Related Work
In audio-driven talking head synthesis, one-shot methods have garnered research attention due to their ability to minimize data requirements to a single photograph. While early GAN-based approaches have succeeded in generating photorealistic textures and accurate lip movements, they fall short in capturing realistic expressions and head movements. Recent diffusion-based methods have made substantial progress in addressing these limitations, truly achieving the generation of realistic talking-head videos without relying on carefully curated templates. EMO [3] stands out as a groundbreaking approach that leverages audio and other “weak controls" to train a conditional diffusion model for end-to-end video generation, enhancing the vividness and realism of the generated results. Subsequent methods have widely adopted its innovative Dual U-Net architecture and two-stage training scheme. Notable follow-up works include EchoMimic [14], Hallo [15], and Loopy [13], which have improved the original framework in various aspects such as hierarchical audio-visual mapping, diverse driving signals, and long-term motion modeling, respectively. Nonetheless, these methods still rely on denoising and rendering images in a general VAE space. This redundant, implicit latent space not only increases the learning complexity of the Diffusion model but also results in slow inference speeds. VASA-1 [17] introduces a novel approach by training a Diffusion Transformer (DiT) [18] within a motion-appearance disentangled space, defined by a face reenactment model [19]. This latent space, tailored specifically for the facial domain, combined with a two-stage generation process, significantly reduces inference time while maintaining vibrant results, which has inspired our work. While, its use of implicit face representation lacks correspondence with facial semantics, hindering regional control and detail refinement in the generated results.
3 Method
Our approach enables the generation of realistic talking-head videos from a single portrait image and arbitrary speech audio clips, meanwhile supporting fine-grained control over facial attributes including emotions, headpose, and gaze. We begin by constructing a motion-appearance disentangled facial latent space, which serves as a bridge between motion generation and photorealistic rendering modules. Subsequently, we design a Diffusion Transformer (DiT) to learn the audio-to-motion mapping within this representation space. Finally, we present inference optimization for realtime deployment.
3.1 Motion Space
While Latent Diffusion Models (LDMs) [20] significantly reduce computational requirements by operating in latent space rather than pixel space, their pretrained VAE-defined latent representations present two key limitations in talking head synthesis. (1) The latent space remains redundant, with entangled motion and texture, introducing overhead in both diffusion training and pixel-level rendering. (2) These representations are implicit and lack explicit correspondence with facial attributes, limiting fine-grained control over the generation process. In the following sections, we demonstrate how our proposed motion space addresses these limitations.
Figure 2: Examples of dedicated fine-grained control of subtle facial motion. The notation “15-0" in the top-left image denotes the application of an x-axis offset to the 15th keypoint. The central image depicts the rendered result when applying zero offsets across all dimensions.
3.1.1 Disentangling Identity and Motion
Due to the shared underlying facial geometry, individuals across different identities exhibit highly consistent facial motion patterns. By constraining the diffusion model to learn only the mapping from audio to universal motions, while incorporating identity information during pixel generation, we can significantly reduce both the training and rendering complexity. The key to achieving this lies in the effective disentanglement of identity and motion. Inspired by recent facial reenactment works [21, 22, 23, 24, 25, 26], we construct an explicit motion space based on expression deformation and head rotation.
Specifically, a single frame $I$ is processed through Motion Extractor $\mathcal{M}$ derived from [27] to yield canonical keypoints $\mathbf{c}\in\mathbb{R}^{K\times 3}$ , expression deformations $\boldsymbol{\delta}\in\mathbb{R}^{K\times 3}$, and headposes $\mathbf{R}\in\mathbb{R}^{3\times 3}$ with translations $\mathbf{t}\in\mathbb{R}^{3}$. The $\mathbf{m}=\{\boldsymbol{\delta},\mathbf{R},\mathbf{t}\}$ serve as the identity-agnostic motion representation, which is used for training the diffusion model to predict the corresponding audio-driven $\mathbf{\hat{m}}=\{\boldsymbol{\hat{\delta}},\mathbf{\hat{R}},\mathbf{\hat{t}}\}$, as illustrated in Figure 1.
Then we devise a one-shot face renderer comprising an Appearance Feature Extractor $\mathcal{F}$, a Face Renderer $\mathcal{G}$ [27] to synthesize talking head videos. Given a target portrait, $\mathcal{M}$ first extracts its canonical keypoints $\mathbf{c_{ref}}$ and motion representation $\mathbf{m_{ref}}=\{\mathbf{\boldsymbol{\delta}_{ref}},\mathbf{R_{ref}},\mathbf{t_{ref}}\}$. Then the reference and generated implicit 3D keypoints ($\mathbf{x_{ref}}$ and $\mathbf{\hat{x}}$) are computed as:
$\mathbf{x_{ref}}$ $=\mathbf{c_{ref}}\mathbf{R_{ref}}+\mathbf{\boldsymbol{\delta}_{ref}}+\mathbf{t_{ref}},$ (1)
$\mathbf{\hat{x}}$ $=\mathbf{c_{ref}}\mathbf{\hat{R}}+\boldsymbol{\hat{\delta}}+\mathbf{\hat{t}}.$ (2)
These two sets of implicit keypoints guide $\mathcal{G}$ in learning a warping field that transforms the appearance features $\mathbf{f_{ref}}$ from $\mathcal{F}$ to fit the generated motions and decodes the warped features into image space to synthesize frame $\hat{I}$ as:
$\hat{I}$ $=\mathcal{G}\left(\mathbf{f_{ref}},\mathbf{x_{ref}},\mathbf{\hat{x}}\right).$ (3)
3.1.2 Motion Control
Unlike headpose which can be directly controlled through rotation and translation, 3D implicit keypoint-based motion representations [21, 22, 23] lack clear correspondence with facial attributes. To address this limitation, [27] employs landmark-guided optimization to align compact implicit keypoints with explicit landmarks. We extend this work by establishing the direct mapping between deformations and facial semantics, enabling motion control in a blendshape-based manner. Specifically, we represent expression deformation as a 63-D vector, corresponding to the x, y, and z coordinates of 21 implicit 3D keypoints. For each dimension, we apply small positive and negative offsets and render facial images for a fixed identity. As illustrated in Figure 2, each dimension of deformations affects relatively distinct facial regions, generating corresponding local variations. For instance, the 34th dimension governs the right eye’s opening and closing motion. The 58th dimension controls mouth open, analogous to the “jaw open" blendshape in ARKit. Therefore, we enable two types of generation control: (1) Regional control, where motion generation can be constrained to specific local facial areas. For example, limiting movements to facial features close to the neck for better integration with the torso, and (2) Magnitude control, where we can impose constraints on deformation values to prevent artifacts such as unnatural facial expressions.
Figure 3: Visualization of how ICS (Initial Conditional Signal) and ECS (Enhanced Conditional Signal) are fed into our DiT-based Motion Generator.
3.2 Motion Generation with Diffusion Transformer
Consistent with prior works [13, 15, 14], we employ diffusion models for audio-conditioned video generation. Differently, we eschew the Dual-Unet structure [3] and two-stage training in both spatial and temporal stages. Instead, we utilize a transformer network [28], leveraging the compact motion representations, to accomplish the motion generation in time series.
3.2.1 Conditional Information
As shown in Figure 3, we employ a DiT for audio-to-motion generation. Along with the audio feature $\mathbf{a}$, we incorporate multiple auxiliary conditional signals as follows. We use the canonical keypoints as the identity feature $\mathbf{c_{ref}}$ to guide the generated facial motions, ensuring compatibility with the target’s facial structure. This identity-adaptive setting mitigates the long-standing issue of incomplete disentanglement between motion and identity [29, 30]. Using a visual emotion classifier [31], we extract emotions from video frames and employ a unique emotion label $\mathbf{s}$ at the clip level, establishing more direct and controllable correspondence compared to traditional audio-to-expression mapping. Eye state $\mathbf{e}$ is composed of the aspect ratio and relative pupil position [32], governing blinking and gaze which have weak correlations with audio. Dedicated conditions are beneficial for generating accurate results. A reference initial motion $\mathbf{m_{ref}}$ serves as a guide for the motion generation within each clip, which can be utilized to enhance inter-clip motion continuity, as well as to periodically steer the motion towards a specified target state, thereby mitigating error accumulation in long sequence generation [33].
For $\mathbf{e}$, $\mathbf{c_{ref}}$, and $\mathbf{s}$, we use them as the Enhanced Conditional Signals (ECS) which are aligned temporally with audio features $\mathbf{a}$, followed by channel-wise concatenation before feeding into a cross-attention module to guide the generation process. As for $\mathbf{m_{ref}}$, we use it as an Initial Conditional Signal (ICS) by replicating and concatenating it channel-wise with the noise sequence, aiming to influence motion generation primarily in the initial stages. In summary, we denote our input condition as $\mathbf{C}=\left\{\mathbf{a},\mathbf{e},\mathbf{c_{ref}},\mathbf{s},\mathbf{m_{ref}}\right\}$
3.2.2 Training Strategy
We devise specialized strategies for training the DiT to accommodate the characteristics of the talking head scenario and the proposed motion representation.
Horizontal Flip: Facial motions associated with speech are predominantly symmetrical. However, in-the-wild talking-head training data often exhibits an uneven distribution of head orientations. This imbalance can affect the extracted motion features, potentially biasing generated motions towards one side. To address this issue, we horizontally flipped the face images, to augment the training data to a balanced audio-to-motion correspondence for both sides.
Adaptive Loss Weights: The correlation between audio and lip movements is highly consistent across different identities. However, the relationship between audio and facial expression is quite complicated, furthermore, the correlation between audio and headpose is highly individualistic. It is hard to train an audio-driven model by uniformly weighing different motion representations. Therefore, we group the motion representations and dynamically adjust the weights of each group throughout the training process. Experiments demonstrate that this dynamic weighting strategy improves both convergence speed and the quality of generated motion.
Validation Metric: Given that lip movements are the primary descriptor of facial motion, we employ a lipsync score [1] as our validation metric. We render the generated motions from the validation videos and use this metric to assist in selecting the preferred model checkpoint. This approach addresses a common training issue in diffusion models, where the loss curve often fails to reflect the generation quality accurately.
3.2.3 Loss Function
Following DDPM [34], a denoising network $\mathcal{T}$ is trained by minimizing the mean square error loss as follows:
$\mathcal{L}_{d}=\mathbb{E}_{t\sim\mathcal{U}[1,T],\mathbf{m}_{0},\mathbf{C}}\left[\left\|\mathbf{m}_{0}-\mathcal{T}(\mathbf{m}_{t},\mathbf{C},t)\right\|_{2}^{2}\right],$ (4)
where $\mathbf{C}$ is the conditional signal, $\mathbf{m}_{t}$ is a perturbed version of real motion data $\mathbf{m}_{0}$ by adding $t$ step noises. To enhance temporal stability, we also regress the velocities and accelerations of motion movements.
$\mathcal{L}_{t}=\left\|\mathbf{\hat{m}^{\prime}}-\mathbf{m}^{\prime}\right\|_{2}^{2}+\left\|\mathbf{\hat{m}^{\prime\prime}}-\mathbf{m}^{\prime\prime}\right\|_{2}^{2},$ (5)
where $\mathbf{m}^{\prime\prime}$ and $\mathbf{m}^{\prime}$ denote the first-order and second-order derivatives of motion representations $\mathbf{m}$. Furthermore, to enhance the initial motion guidance, we compute an initial motion loss $\mathcal{L}_{ini}$ between the reference initial motion and the first element of the generated motion clip. In summary, the final loss can be expressed as:
$\mathcal{L}=\mathcal{L}_{d}+\mathcal{L}_{t}+\mathcal{L}_{ini}$ (6)
      Method       FID $\downarrow$       FVD $\downarrow$       CSIM $\uparrow$       Sync-C $\uparrow$       Sync-D $\downarrow$       RTF $\downarrow$
      GT       -       -       -       8.044       6.943       -
      MuseTalk       21.445       436.862       0.807       5.586       8.400       2.248
      EchoMimic       42.554       395.754       0.840       5.733       9.204       35.528
      Hallo       22.996       271.680       0.812       7.652       7.590       53.082
      Hallo2       22.899       245.236       0.806       7.737       7.608       56.838
      Ours-s50       17.254       219.368       0.864       8.069       7.114       2.121
      Ours-s10       17.060       231.182       0.861       8.111       7.291       0.635
Table 1: Comparisons with existing methods. The s* represents the number of denoising steps. The best and second best results are in bold and underline specifically.
3.3 Realtime Streaming Inference
Realtime dialogue scenarios impose two critical requirements on talking head methods: (1) low-latency inference within each module and (2) efficient data streaming between modules. To cope with these requirements, we optimize the inference process across three primary modules: audio feature extraction, motion generation, and video synthesis.
3.3.1 Audio Feature
We employ the HuBERT [35] for audio feature extraction, incorporating transformer [36] acceleration techniques to enable realtime streaming audio processing. Specifically, we adopt a KV cache strategy [37] to provide pseudo-contextual information extracted from a fixed corpus, concatenated with incoming audio stream segments. This approach allows the HuBERT model pre-trained on long utterances to maintain high-quality features even on extremely short audio segments. Complementarily, we implement causal mask to ensure proper attention allocation to the actual input while significantly reducing computational complexity. These optimizations enable our model to process 0.4-second audio stream units in realtime on CPU environment. We utilize these enhanced features for both model training and inference.
Figure 4: The qualitative comparison of our approach on two characters with different styles, poses, and hairstyles. For each character, we generat videos using the same audio input with each method and select frames at the same location for comparison. Blue arrows indicate the locations of artifacts, including inaccurate lip movements, disordered teeth, blur, etc.
3.3.2 Motion Generation
We align the audio features with the video frame rate and segment the streaming audio features into fixed-length intervals, incorporating a predefined overlap between adjacent segments. We leverage the long video generation strategy introduced in [38], where we substitute progressive latent fusion with segment-wise fusion on generated motion sequences to support streaming output. During fusion, weights are determined by the relative position of each frame within its segment, with frames closer to the segment center receiving higher weights.
To achieve realtime performance, we reduce the number of denoising steps in DiT inference from 50 to 10. Benefiting from the motion representation, the inference results at 10 and 50 steps maintain comparable quality in evaluation. The DiT model is converted to TensorRT for execution on GPU environment.
3.3.3 Video Synthesis
Our one-shot face renderer sequentially generates photorealistic images from motion sequences. Features relevant to the target identity are pre-extracted and stored. The TensorRT-optimized renderer enables realtime inference and video stream output on GPUs. If video output is required, a CPU-based FFmpeg can parallelly compress the video stream.
Figure 5: Generation results with portraits of different styles and scales. Fine-grained control over gaze, emotion, pose, etc.
4 Experiments
4.1 Implementation Details
For training data, we collect and clean about 50 hours of broadcast scene videos from 330 identities, with an average video length of 150 seconds. For testing, we use the videos collected by previous works [39, 40, 41, 42] for quantitative evaluations. To evaluate the generalizability of our method, we select diverse image and audio samples from [15, 14] for qualitative evaluations.
In experiments, the headpose euler angles are represented using discrete bins [21], and the final motion representation contains 265 dimensions. We have trained our models using 8 NVIDIA A100 GPUs with a batch size of 1024, using the Adan [43] optimizer with a learning rate of 1e-4 and 0.02 weight decay for 500 epochs. We select the preferred checkpoint based on the lipsync score.
During training, audio features, video frames, and corresponding conditional signals are aligned at 25 fps. We sample segments of $L$=80-frames (3.2s) from video clips to train the Diffusion Transformer. The reference initial motion $\mathbf{m_{ref}}$ is derived from the preceding frame of the current training segment, and the reference canonical keypoints $\mathbf{c_{ref}}$ are taken from a random frame of the same video clip. During inference, both the $\mathbf{m_{ref}}$ and $\mathbf{c_{ref}}$ are derived from the source image.
4.2 Baseline and Metrics
We compare our method with recent publicly available implementations, including the non-diffusion-based lip editing method MuseTalk [44], and the diffusion-based talking head generation methods EchoMimic [14], Hallo [15] and Hallo2 [45]. The following metrics are used to evaluate the generated results: Fréchet Inception Distance (FID) [46] for frame-level quality, Fréchet Video Distance (FVD) [47] for video-level quality, Cosine Similarity (CSIM) of face recognition features [48] for identity consistency, Sync-C and Sync-D [1] for audio-visual synchronization. We also evaluate the inference performance by Real-Time Factor (RTF) and First-Frame Delay (FFD). $\text{RTF}<1$ and an FFD as low as possible are the basic requirements for realtime streaming inference.
4.3 Quantitative Evaluation
The quantitative results are shown in Table 1. Our method achieves the best results among competing methods on all evaluated metrics. The compact motion representation facilitates the learning of the diffusion model and makes the lipsync score close to the real video, which is the most important criterion for evaluating talking head tasks. How to maintain identity consistency, measured by the CSIM score, is a long-standing challenge for one-shot methods. Our method achieves higher CSIM scores than other methods, attributed to the disentanglement of appearance and motion. Our method also effectively improves the frame quality and inter-frame continuity of the generated video obtaining the lowest FID and FVD among all methods. Table 1 also reports the results of reducing the number of denoising steps. The results of 10-step and 50-step denoising exhibit minimal disparity in end-to-end evaluation, substantiating the efficacy of our inference optimization. Analysis of the RTF reveals that diffusion-based methods exhibit inference speeds 30-50 times slower than realtime. In contrast, our proposed method achieves inference capabilities that surpass realtime performance, and even outpace the non-diffusion-based lip-editing method MuseTalk.
Method FID $\downarrow$ FVD $\downarrow$ CSIM $\uparrow$ SyncC $\uparrow$ SyncD $\downarrow$
Full Model 17.254 219.368 0.864 8.069 7.114
w/o C-kp 17.438 261.693 0.849 6.932 8.476
w/o Emo 23.036 324.078 0.792 6.310 8.990
w/o Ada-w 18.978 588.821 0.844 0.289 14.992
Table 2: Ablation study results. C-kp represents canonical keypoints condition, Emo represents emotion label condition, and Ada-w represents Adaptive Loss Weights.
4.4 Qualitative Evaluation
Figure 4 shows a qualitative comparison of our approach on two characters with different styles, poses, and hairstyles. For each character, we generate videos using the same audio input with each method and selected frames at the same location for comparison. Our method shows a notable edge in the consistency of generated details. For example, the shape, number and texture of teeth generated by other methods may change between frames, while our generation results show consistency between consecutive frames. The comparison methods also have problems such as generating exaggerated expressions, blur, and inaccurate lip movements. These problems are unacceptable concerns to customers in real application scenarios. In addition, the generation results of Hallo and Hallo2 for stylized portraits may deteriorate gradually over time, such as excessive facial wrinkles, abnormal shadows, and background artifacts. In contrast, our method demonstrates consistent performance across diverse portrait styles.
Figure 5 demonstrates that Ditto can not only handle portraits of different styles but also support fine-grained generation control, such as gaze, pose, emotion, etc. Furthermore, our approach enables seamless stitching of the generated head with the original body, facilitating full-body portrait animation without the misalignment issues often associated with end-to-end diffusion methods. This makes Ditto readily integrate with the body-driven model, achieving strong generation capabilities in a wide range of applications.
4.5 Ablation Study
Conditional Information. We evaluate the impact of different controlling signals on the end-to-end generation results. As shown in Table 2, when canonical keypoints are excluded (w/o C-kp), all metrics drop, especially for FVD and Sync. This suggests that the identity-related geometry provided by canonical keypoints enhances the learning process of diffusion, particularly in adapting generated motion to the target identity. Conversely, imperfections in motion generation directly affect dynamic visual quality and lip synchronization accuracy.
When emotion labels are omitted (w/o Emo), we observe a significant decline across all metrics. Without direct emotional control, learning facial expressions solely from audio becomes more challenging, resulting in a decrease in video quality and lip synchronization. In the supplementary materials, we demonstrate the consequent issues such as abrupt expression changes and exaggerated facial expressions, which also impact identity consistency throughout the video.
Adaptive Loss Weights. Without adaptive loss weights (w/o Ada-w), the problem of unbalanced convergence between pose and expression becomes severe, hindering the model’s ability to learn the correspondence between audio and facial motion. We observe that the generated videos exhibit minimal facial motions, resulting in poor Sync and FVD scores, while CSIM and FID metrics are less affected.
4.6 Inference Performance
Module Audio2Feat Motion DiT Face Rendering
inference 23ms 62ms 15ms
RTF 0.115 0.310 0.375
Table 3: The time consumption of single-step inference in different modules and corresponding real-time factor (RTF).
     Pipeline      RTF      FFD
     talking-head offline      0.635      -
     talking-head online      0.895      385ms
     full-body offline      0.648      -
     full-body online      0.914      392ms
Table 4: Inference performance for offline video output and online streaming output at the head region (512 $\times$ 512) and full-body scales (1920 $\times$ 1080).
We separately evaluate the inference performance of the proposed method from both modular and end-to-end perspectives. We evaluate module-wise performance using both inference time and valid Real-Time Factor (RTF), which represents the single-step inference time divided by the valid segment length (excluding the overlapping part). The test environment is a 12-core Intel(R) Xeon(R) Platinum 8369B CPU @ 2.90GHz, 1 NVIDIA A100 GPU, and 100G memory.
Table 3 shows that each module’s RTF is under 1, guaranteeing realtime end-to-end inference for the pipeline. In Table 4, we present the inference performance for offline video output and online streaming output at two scales: the head region and full-body portrait. Compared to offline video output, online streaming output requires longer inter-segment overlap to ensure low First-Frame Delay (FFD), resulting in a higher RTF. Full-body output is only marginally slower than head-only output, primarily due to the process of integrating the head into the body. Our method ensures end-to-end inference with RTF below 1 and FFD under 400ms, meeting the requirements for realtime interactive scenarios.
5 Conclusion
We propose a diffusion-based framework specifically designed for one-shot talking head synthesis. We use a Diffusion Transformer to generate motion sequences in an identity-independent space and synthesize videos through a dedicated face renderer. It can produce realistic videos from a single portrait image and support fine-grained motion control, promising for realtime interactive applications.